The Admissions officer for the graduate programs at the University of Pennsylvania believes that the average score on the LSAT exam at his university is significantly higher than the national average of 1300. An accepted standard deviation for LSAT scores is 125. A random sample of 25 scores had an average of 1375.
c) Calculate the p-value.

Answer by stanbon(75887)

The Admissions officer for the graduate programs at the University of Pennsylvania believes that the average score on the LSAT exam at his university is significantly higher than the national average of 1300. An accepted standard deviation for LSAT scores is 125. A random sample of 25 scores had an average of 1375.
Ho: u <= 1300
Ha: u > 1300 (claim)
---------------------
t(1375) = (1375-1300)/[125/sqrt(25)] = 75/25 = 3
-----
c) Calculate the p-value::
p-value = P(t > 3 when df = 24) = tcdf(3,100,24) = 0.0031
-----------------
Cheers,
Stan H.
----------------